On May 1, 2021, Hecala Mining entered into an agreement with the state of New Mexico to obtain the rights to operate a mineral mine in New Mexico for $11.0 million. Additional costs and purchases included the following (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.): Development costs in preparing the mine $ 4,200,000 Mining equipment 146,000 Construction of various structures on site 121,500 After the minerals are removed from the mine, the equipment will be sold for an estimated residual value of $11,000. The structures will be torn down. Geologists estimate that 900,000 tons of ore can be extracted from the mine. After the ore is removed the land will revert back to the state of New Mexico. The contract with the state requires Hecala to restore the land to its original condition after mining operations are completed in approximately four years. Management has provided the following possible outflows for the restoration costs: Cash Outflow Probability $ 700,000 30% 800,000 40% 900,000 30% Hecala’s credit-adjusted risk-free interest rate is 9%. During 2021, Hecala extracted 130,000 tons of ore from the mine. The company’s fiscal year ends on December 31. Required: 1. Determine the amount at which Hecala will record the mine. 2. Calculate the depletion of the mine and the depreciation of the mining facilities and equipment for 2021, assuming that Hecala uses the units-of-production method for both depreciation and depletion. 3. How much accretion expense will the company record in its income statement for the 2021 fiscal year? 4. Are depletion of the mine and depreciation of the mining facilities and equipment reported as separate expenses in the income statement? 5. During 2022, Hecala changed its estimate of the total amount of ore originally in the mine from 900,000 to 1,100,000 tons. Calculate the depletion of the mine and depreciation of the mining facilities and equipment for 2022 assuming Hecala extracted 160,000 tons of ore in 2022.
Answer #1

Solution

Hecala Mining

  1. Determination of the amount at which Hecala will record the mine:

Mining site cost

$11,000,000

Restoration costs -

Development costs

$4,200,000

[(700,000 x 30%) + (800,000 x 40%) + (900,000 x 30%)] x (P/F, 9%, 4)

Restoration Costs

$566,720

$800,000 x (P/F, 9%, 4) = 800,000 x 0.7084 = $566,720

Cost of the mine

$15,766,720

  1. Calculation of the depletion of the mine and depreciation fo the mining facilities and equipment for 2021 assuming units of production method for both depreciation and depletion:

Depletion = cost of mine per ton x tons extracted

= ($15,766,720/900,000 tons) x 130,000 tons

Depletion = 2,277,415

Depreciation – Machinery

Depreciation = (depreciable base/estimated tons) x tons extracted

Depreciable base = cost – residual value

= 146,000 – 11,000 = $135,000

Depreciation rate = $135,000/900,000 tons = $0.15 per ton

Depreciation expense = $0.15 x 130,000 tons = $19,500

Hence, machinery depreciation expense = $19,500

Depreciation – Structures

Depreciation = (cost/estimated tons) x tons extracted

= ($121,500/900,000 tons) x 130,000 tons

= $17,550

Hence, depreciation expense – structures = $17,550

  1. Determination of the accretion expense to be recorded in the income statement for the 2021 fiscal year:

Accretion expense = restoration cost x risk free interest rate x 8/12 months

= $800,000 x 9% x 8/12 = $48,000

  1. The depletion of mine and depreciation of mining facilities, equipment and structure form part of income statement as cost of goods sold only when the company sells the mineral.
  2. Computations of depletion and depreciation with change in estimates for 2022 –

Estimated mine = 1,100,000 tons; extracted mine = 160,000 tons

Depletion –

Cost of mine = $15,766,720

Less: depletion 2021 = $2,277,415

Book value at start of 2022 = $13,489,305

Revised estimates of tons remaining = 1,100,000 – 130,000 = 970,000 tons

Depletion rate = $13,489,305/970,000 = $13.9065 per ton

Tons extracted in 2022 = 160,000

Depletion 2022 = $13.9065 x 160,000 = $2,225,040

Depletion for 2022 = $2,225,040

Depreciation of machinery –

Book value at start of 2022 = depreciable base – depreciation 2021

= $135,000 - $19,500 = $115,500

Revised estimate of tons remaining = 970,000

Depreciation rate = $115,500/970,000 = $0.11907

Depreciation expense = 160,000 x $0.11907 = $19,052

Machinery Depreciation expense, 2022 = $19,052

Depreciation of structures –

Book value at start of 2022 = $121,500 – 17,550 = $103,950

Revised estimate of tons remaining = 970,000 tons

Depreciation rate = $103,950/970,000 = $0.10716

Depreciation expense 2022 = $0.10716 x 160,000 tons = $17,146

Structures Depreciation expense 2022 = $17,146